OpenStudy (anonymous):

One of the qualifications for a job is a typing speed of at least 65 words per minute. If Jordan knows that she will be able to type 350 words in five minutes, what is the maximum number of errors she can make? Select one: A.) 4 B.) 0 C.) 3 D.) 2

OpenStudy (anonymous):

so what you do is you divide 350 by five to see how may words per minute she types then you subtract 65 from the answer you get to show how many she can make.

OpenStudy (anonymous):

but then ur answer is 5 and non of the selections are 5 @theluckyistduck

OpenStudy (anonymous):

Hm idk but it makes sense to see how many words per minute she types and the subtract 65 from her words per minute

OpenStudy (anonymous):

ik and i thought of the samething as u but non of the answers r 5 so i think we r both wrong and r missing something.

OpenStudy (anonymous):

hmm @Ashley1998 are you sure none of the answers are 5?

OpenStudy (anonymous):

they are 4. 0. 3. and 2.

OpenStudy (agent0smith):

This is an odd question... in five mins she needs to type 5x65 = 325 words. She can type 350 words in 5 minutes, so she can make mistakes on 25 words total, or 5 mistakes per minute...

OpenStudy (anonymous):

yea that's my reasoning to this question too..

OpenStudy (anonymous):

could it be 350 / 5 = 70 - 66?

OpenStudy (anonymous):

maybe it is a trick question and the answer is 0 cause the job wants her to make no mistakes :P

OpenStudy (anonymous):

y 66 Ashley?

OpenStudy (anonymous):

70-66=4

OpenStudy (anonymous):

lol yes but where did u get 66?

OpenStudy (anonymous):

Idk but adding 66 in doesn't make sense.

OpenStudy (kropot72):

A Word = 5 characters, including spaces. Time is measured in Elapsed Seconds [ES ]. The formula: Adjusted WPM = { Words − Errors } x [60/ES]

OpenStudy (anonymous):

cause if it is to get one of those answers then u could of done either of these solutions 70-66 70-67 70-68 70-69 70-70

OpenStudy (anonymous):

oh wow, there is a formula for this too?!?! lol they make those for everything xD

OpenStudy (anonymous):

I know, but the do come in handy

OpenStudy (anonymous):

i think he made that one up >.>

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I still say the maximum amout of errors she can make is 5. I would suggest you ask your teacher about it.

OpenStudy (anonymous):

well as nice as the formula is, i dont know how to use it :/

OpenStudy (anonymous):

*amount ya I tried to use it and it didn't make sense.

OpenStudy (anonymous):

@theluckyistduck accept ur wrong ! :P

OpenStudy (anonymous):

infinity sign guy, help a lil more

OpenStudy (anonymous):

I know but is the only way that solved out makes sense!!!!!

OpenStudy (anonymous):

i sorry guys i didnt see but here is this For exercises 4-5 use this equation to estimate typing speed S= W/5 – 2e, where S is the accurate typing speed, W is the number of words typed in 5 minutes, and E is the number of errors.

OpenStudy (agent0smith):

Ah, there you go Ashley :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

lol im sorry

OpenStudy (anonymous):

i feel less stupid now! :D

OpenStudy (anonymous):

ok gl with the question!

OpenStudy (agent0smith):

S= W/5 – 2e so you need to solve this for e. W = 350, i think S has to be 60.

OpenStudy (anonymous):

no e=70, cause 350words = 5mins 70words = 1 min

OpenStudy (anonymous):

speed is distance PER unit time

OpenStudy (anonymous):

and the answer was 0 :P so i was right all along :D

OpenStudy (anonymous):

haha good for you!!!

OpenStudy (anonymous):

see so this is a trick question :P i saw straight through lol and thx @theluckyistduck !

OpenStudy (anonymous):

lol thank you guys for helping me :D

OpenStudy (anonymous):

yes yes you did. Sorry it took so long

OpenStudy (anonymous):

well we didnt quite help >.>

OpenStudy (agent0smith):

@HawkCrimson isn't e= errors not words per minute?

OpenStudy (anonymous):

i was talking about S

OpenStudy (anonymous):

lol its cool, i take all the time i need cuz i suck at math! lol

OpenStudy (anonymous):

S= W/5 – 2e 70 = 350/5 - 2e 70 = 70 - 2e -2e = 0 e = 0

OpenStudy (agent0smith):

S= W/5 – 2e W = 350, i think S has to be 60. 60 = 350/5 - 2e

OpenStudy (agent0smith):

S is the required accurate typing speed.

OpenStudy (anonymous):

dude, ur answer is 5 so ur wrong

OpenStudy (anonymous):

so then where did the 60 come from? Isn't it 65?

OpenStudy (agent0smith):

Oh, yes, 65.

OpenStudy (agent0smith):

65 = 350/5 - 2e

OpenStudy (anonymous):

no, 70 -_-

OpenStudy (agent0smith):

65 is the accurate required typing speed.

OpenStudy (anonymous):

350 words = 5mins 70 words = 1 min therefore, 70words per min, which is her typing speed

OpenStudy (agent0smith):

e is not zero...

OpenStudy (agent0smith):

again, S is the required accurate typing speed.

OpenStudy (anonymous):

well it is either 0 or 2.5 so I would take the answer that is listed in the possible answers over the one that is not.

OpenStudy (anonymous):

S= W/5 – 2e 65 = 350/5 - 2e 65 = 70 - 2e -5 = -2e e = 5/2

OpenStudy (agent0smith):

use this equation to estimate typing speed S= W/5 – 2e, where S is the accurate typing speed, W is the number of words typed in 5 minutes, and E is the number of errors. Hence S = 65 and e = 2.5, which means max errors is really 2.

OpenStudy (anonymous):

does any of the answer look like 5/2 to u????? then im right :P

OpenStudy (agent0smith):

You are not right... all you did was put in the same value for S as W/350.

OpenStudy (anonymous):

dude ur wrong, accept it

OpenStudy (anonymous):

it is ok man everyone gets it wrong from time to time :P

OpenStudy (agent0smith):

S is not W/350, S is required accurate typing speed. W is words typed per minute. Please read: use this equation to estimate typing speed S= W/5 – 2e, where S is the accurate typing speed, W is the number of words typed in 5 minutes, and E is the number of errors.

OpenStudy (anonymous):

infinity guy will settle it :D

OpenStudy (agent0smith):

S is the accurate typing speed, W is the number of words typed in 5 minutes, and E is the number of errors. in other words, S must be 65. W is .... 350!

OpenStudy (agent0smith):

e = 2.5, and thus she can make max 2 errors, since 3 will push her over that required limit.

OpenStudy (anonymous):

S is the number of ACCURATE, not aquired

OpenStudy (anonymous):

and her accurate typing speed is 70 words per min

OpenStudy (agent0smith):

No, her typing speed is 70 words per minute. That isn't necessarily her accurate typing speed.

OpenStudy (anonymous):

i garantee that it is more accurate than 65

OpenStudy (kropot72):

There must be an allowable number of errors on condition the typist can exceed the specified typing speed. The typist in the question can exceed the specified typing speed therefore the allowable number of errors cannot be zero in this case. Using the formula given with the original question I get maximum errors = 2.5 which must be rounded down to 2 errors.

OpenStudy (anonymous):

65 is the general accepted speed and 70 is her accurate speed

OpenStudy (agent0smith):

...the question is asking how many errors she can make. If she makes zero errors... what's the point of the question?

OpenStudy (agent0smith):

thank you @kropot72

OpenStudy (kropot72):

@agent0smith You're welcome :)

OpenStudy (anonymous):

no ur making a false assumption

OpenStudy (agent0smith):

@HawkCrimson if she could accurately type 70 words per minute with zero errors, we wouldn't be discussing the max number of errors she can make. It *never* states she can *accurately* type 70 words per minute, only that she CAN type 70 a minute.

OpenStudy (anonymous):

this is not a real life senario, in RL yes but for the sake of this question ur making an assumption that cannt nessassarily be true

OpenStudy (agent0smith):

...what?

OpenStudy (anonymous):

and her max is 0 :P

OpenStudy (agent0smith):

Her max would be zero if she could only type 65 words per minute...

OpenStudy (anonymous):

and they never said that she couldnt accuratly type 70 so u cannt assume other wise

OpenStudy (agent0smith):

doesn't that seem logical? If she can type 65 words a minute, she can't afford a single error, right?

OpenStudy (agent0smith):

If she can type more than 65 a minute... she can afford mistakes.

OpenStudy (agent0smith):

She may be able to type 70 a minute accurately... thats not the question. it's what is the max number of errors she CAN make? Not will make.

OpenStudy (anonymous):

u people r just making assumptions but it doesnt nessassarily need to be true

OpenStudy (anonymous):

oh ok i get it now :P

OpenStudy (agent0smith):

If she can type over 65 words a minute, she can make mistakes, right?

OpenStudy (anonymous):

u see this is how i learn fro giving my teachers hell :D

OpenStudy (agent0smith):

haha :)

OpenStudy (anonymous):

so the minimum mistakes she can make is 0, correct?

OpenStudy (anonymous):

we agree on that much?

OpenStudy (agent0smith):

Well yeah, she can't make negative mistakes :P

OpenStudy (anonymous):

:D

OpenStudy (agent0smith):

minimum has to be zero :P

OpenStudy (anonymous):

lol as long as something is equal to 0 im happy :P